# Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation

### v1

@@ 2.2 Theoretical conditions for identifiability and inference
Lemma 2.2.
The bias ${\mathbf{b}}$ can be effectively utilized for prediction if Assumptions A1 through A5 hold.
As shown in Lemma 2.1, under Assumptions A1–A4, both content and bias information ${\mathbf{c}},{\mathbf{b}}$ can be well identified. It indicates that we can extract ${\mathbf{c}},{\mathbf{b}}$ containing all the information as the ground-truth variables. Assumption A5 states that
$P(y\mid{\mathbf{c}},{\mathbf{b}})\;\neq\;P(y\mid{\mathbf{c}})\quad\text{on a set of nonzero measure.}$.
In Bayesian decision theory (for any proper loss, e.g. log-loss or squared error), the Bayes-optimal predictor using $({\mathbf{c}},{\mathbf{b}})$ is
$f^{*}({\mathbf{c}},{\mathbf{b}})=\arg\min_{f}\;\mathbb{E}\bigl{[}\ell\bigl{(}y,f({\mathbf{c}},{\mathbf{b}})\bigr{)}\bigr{]},$
which depends on the full conditional $P(y\mid{\mathbf{c}},{\mathbf{b}})$. The best predictor using only ${\mathbf{c}}$ is
$\arg\min_{f}\;\mathbb{E}\bigl{[}\ell(y,f({\mathbf{c}}))\bigr{]},$
which depends only on $P(y\mid{\mathbf{c}})$. Since Assumption A5 ensures that the conditional distributions $p(y\mid{\mathbf{c}})$ and $p(y\mid{\mathbf{c}},{\mathbf{b}})$ differ, the joint predictor $f^{*}({\mathbf{c}},{\mathbf{b}})$ achieves a strictly lower expected loss compared to any predictor that relies solely on ${\mathbf{c}}$. This confirms that the bias variable ${\mathbf{b}}$ can indeed be exploited to improve predictive accuracy. A detailed proof, demonstrating that incorporating ${\mathbf{b}}$ leads to a better predictor under log-loss or squared error, is provided in Appendix B.2.
2.3 Understand bias’ contribution in prediction
Furthermore, we explore how bias contributes to the predictions. To illustrate the core insights of our approach, we begin with a binary classification setting. The extension to the multi-class scenario is in Appendix C. The core intuition is that $\mathbf{c}$ and $\mathbf{b}$ are conditionally independent given $y$, which indicates that we can decompose $p(y\mid\mathbf{c},\mathbf{b})$ into three parts, including a bias-aware predictor, an invariant predictor, and a label prior.
Theorem 2.3.
[Decomposition of Marginal Probability] Given the data generation process illustrated in Figure 2, we have:
$\ p(y=1\mid\mathbf{c},\mathbf{b})=\sigma\biggl{(}\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{b})\bigr{)}}_{\text{Bias-aware Predictor}}+\underbrace{\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c})\bigr{)}}_{\text{Invariant Predictor}}-\underbrace{\operatorname{logit}\bigl{(}p(y=1)\bigr{)}}_{\text{Label Prior}}\biggr{)},$
where $\sigma$ is the sigmoid function. The last two terms combined can be viewed as an invariant part of the model across all domains.
Proof sketch of Theorem 2.3
We begin by reformulating $p(y\mid\mathbf{c},\mathbf{b})$ to make its structure more explicit. Specifically, we have:
$\operatorname{logit}\bigl{(}p(y=1\mid\mathbf{c},\mathbf{b})\bigr{)}=\frac{p(\mathbf{c},\mathbf{b}\mid y=1)\,p(y=1)}{p(\mathbf{c},\mathbf{b}\mid y=0)\,p(y=0)}\quad\quad\text{consider }\mathbf{c}\,\perp\!\!\!\perp\,(\mathbf{b})\,\mid\,y:$
$\quad=\frac{p(\mathbf{b}\mid y=1)\,p(\mathbf{c}\mid y=1)\,p(y=1)}{p(\mathbf{b}\mid y=0)\,p(\mathbf{c}\mid y=0)\,p(y=0)}=\frac{p\bigl{(}y=1\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=1\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=1\bigr{)}}}{p\bigl{(}y=0\mid\mathbf{b}\bigr{)}\,\tfrac{p\bigl{(}y=0\mid\mathbf{c}\bigr{)}}{p\bigl{(}y=0\bigr{)}}}.$
A more detailed proof is provided in Appendix D.1.
Discussion on invariant prediction.
We observe that $p(\mathbf{c}\mid y)$ remains invariant across domains, since $y$ d-separate ${\mathbf{c}}$ and the environment variable $e$. Formally, $\forall\,e_{i},e_{j}\in E_{all}$, since $\mathbf{c}$ only depends on $y$, it follows that
$p(\mathbf{c}\mid y=1,\,e=e_{i})\;=\;p(\mathbf{c}\mid y=1,\,e=e_{j})\Rightarrow\frac{p(y=1,\,e=e_{i}\mid\mathbf{c})}{p(y=1,e=e_{i})}\;=\;\frac{p(y=1,\,e=e_{j}\mid\mathbf{c})}{p(y=1,e=e_{j})}$
which implies $\tfrac{p(y\mid\mathbf{c})}{p(y)}$ is an invariant quantity across all domains.
2.4 Leverage bias for better prediction
As shown in Theorem 2.3, we have established that the bias affects prediction solely through the bias-dependent component $p(y=1\mid\mathbf{b})$. Consequently, ensuring the reliability of this bias-aware prediction becomes crucial for effectively leveraging bias. However, relying on this term without additional assumptions poses challenges, as the bias-aware prediction component can vary significantly across environments. In the absence of proper constraints, incorporating it naively would revert the model to standard empirical risk minimization (ERM), failing to account for bias robustness. Furthermore, as shown in the causal graph in Figure 2, the term $p(y=1\mid\mathbf{b})$ can be decomposed into two components: a direct path $\mathbf{b}\rightarrow y$ and an indirect path $\mathbf{b}\leftarrow e\rightarrow y$. Motivated by this structure, we propose to leverage bias for prediction in both direct and indirect ways. The direct approach uses invariant features to guide the extraction of the predictive (i.e., positively correlated) component of the bias. In contrast, the indirect approach utilizes the bias to infer the environment $e$, which in turn informs the selection of an appropriate predictor tailored to that environment.
2.4.1 Using bias directly
The core idea of directly leveraging bias for the prediction is to extract its beneficial components, guided by the information encoded in invariant features.
Assumption 2.4 (Sufficient Information of ${\mathbf{c}}$).
@@ 2.4.1 Using bias directly
Theorem 2.5 (Correction with Pseudo Labels).
Consider the predictor $f_{c}:{\mathbf{c}}\rightarrow y$. Define pseudo-label prediction as $\hat{y}=f_{c}({\mathbf{c}})$. Let $h_{0}\;=\;p(\hat{y}=0\mid y=0),\quad h_{1}\;=\;p(\hat{y}=1\mid y=1).$ Then it holds that
@@
Justification about Theorem 2.5.
Equation 3 depends and only depends on the 2.4. The detailed proof can be found in Appendix D.2. Intuitively, if $c$ helps distinguish between $y=0$ and $y=1$, then a predictor based on $c$ should be corrected more often than chance, implying $h_{0}+h_{1}>1$.
Proof of Equation 4.

### v2

@@ 5.2 Real-world Experiments
Table 2: Domain-specific OOD accuracy on DomainNet. All methods use ResNet-50 as backbones.
Algorithm Clipart Infograph Painting Quickdraw Real Sketch Avg ERM 58.1 18.8 46.7 12.2 59.6 49.8 40.9 IRM(Arjovsky et al., 2019) 48.5 15.0 38.3 10.9 48.2 42.3 33.9 ACTIR (Jiang and Veitch, 2022) 50.0 22.6 43.5 11.7 57.8 46.8 38.7 RDM (Nguyen et al., 2024) 62.1 20.7 49.2 14.1 63.0 51.4 43.4 GMDG (Tan et al., 2024) 63.4 22.4 51.4 13.4 64.4 52.4 44.6 SFB (Eastwood et al., 2023) 60.7 25.3 49.0 12.0 64.0 50.2 43.5 BAG 61.8 25.6 51.8 13.9 65.5 50.3 44.8
Results and discussions.
For PACS (left in Table 1), our method achieves the best average accuracy, improving by $1.9\%$ over the strongest baseline, GMDG. Compared with the most closely related method, SFB, our model consistently improves performance across all target domains. The largest gain is observed on Art Painting, possibly because this domain contains stronger style/background shifts, where modeling bias is particularly beneficial. On Office-Home (right in Table 1), BAG outperforms all baselines on average. In particular, it achieves improvement on three of the domains, and obtains comparable results on Arts. The notably larger improvement on Clipart which is generally regarded as more challenging, suggests that BAG generalizes better under larger domain shifts. On DomainNet, BAG achieves the best average accuracy, surpassing GMDG and RDM. It attains the best performance on infograph, painting and real, while remaining competitive on clipart , quickdraw, and sketch. The largest gain appears on infograph, where background and stylistic artifacts are prominent, indicating that disentangling content from bias helps suppress spurious domain cues and better exploit informative bias factors. We also outperform ACTIR and SFB by a substantial margin, which we attribute to more accurate disentanglement and a more principled use of bias.
5.3 More Ablation Studies, Sensitivity Check, and Visualization.
Since BAG involves a two-stage optimization procedure and a VAE component, we conduct some ablation studies to quantify the contributions of each design choice. We remove the corresponding components, such as VAE, test-time adaptation that uses pseudo-labels to guide the bias branch, and the reweighting-of-experts module while keeping the rest unchanged as ablation studies. Full ablation details are provided in Section E.2. To verify the stability of BAG, we study its sensitivity to the number of domain experts and the embedding dimensionality in Section E.3. We also visualize the learned representations to illustrate the disentanglement of latent variables in Section E.4.
6 Conclusion
In this paper, we challenge the conventional methods for OOD generalization that treat bias solely as an obstacle. Instead, we theoretically demonstrate that bias can be strategically leveraged to enhance prediction, particularly when it retains useful dependencies on the target domain labels. Our framework demonstrates that bias can be utilized in two ways: bias-aware prediction with an environment routing mechanism and bias correction with an adaptive label prior. Through empirical validation on a synthetic task and some real-world benchmarks, we establish the effectiveness of the proposed approach. The results consistently show that leveraging bias in a structured manner leads to more robust and adaptable models, offering a new perspective on OOD generalization.
Limitations. This paper provides a principled general framework that leverages bias for domain generalization. From a theoretical perspective, the success of such a framework relies on assumptions about the disentanglement of bias features and invariant features. Although the disentanglement is supported by both real-world and simulated experiments, it typically depends on intrinsic structural properties of the data (e.g., linear independence) that are difficult to verify in practice. We therefore recommend users to conduct independence tests between bias and invariant features or the visualizations on the feature space as a practical check metric to help assess their distinguishability in real applications.
Impact Statement
We propose a method that strengthens unsupervised domain generalization, enabling models to transfer knowledge more reliably between different data domains. The approach is designed to improve learning efficiency when annotations are limited or unavailable, which is common in real world deployments. The technique is methodological and does not create new ethical concerns beyond those already associated with standard machine learning systems. Overall, our contribution supports more robust and practical domain generalizing models.
Appendix A Related Work
Domain Generalization Domain Generalization (DG) aims to develop models that generalize to unseen OOD target domains using only observed source domains for training (Zou et al., 2024; Zhang et al., 2024; Chen et al., 2023; Prashant et al., 2024). A common approach is learning invariant representations across all source domains, as seen in Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) and its extensions (Ahuja et al., 2020; Rosenfeld et al., 2020; Ahuja et al., 2021; Kamath et al., 2021; Krueger et al., 2021). Other methods leverage adversarial learning (Li et al., 2018a; Shao et al., 2019; Zhao et al., 2020; Deng et al., 2020; Matsuura and Harada, 2020) to extract domain-invariant features or use disentangled representation learning (Ilse et al., 2020; Wang et al., 2020b; Jiang and Veitch, 2022) to separate domain-invariant from domain-specific information. Causal inference techniques (Wang et al., 2020c; Qi et al., 2020), such as back-door and front-door rules, help mitigate bias. Furthermore, techniques such as data augmentation (Volpi and Murino, 2019; Shi et al., 2020; Huang et al., 2021; Li et al., 2021), feature augmentation (Mancini et al., 2020; Shu et al., 2021; Zhou et al., 2024), and network regularization (Seo et al., 2020; Segu et al., 2023) can reduce overfitting and improve generalization.
Causal Representation Learning Causal Representation Learning (CRL) aims to recover latent variables that correspond to the causal generative factors of high-dimensional observations and the mechanisms that link them (Schölkopf et al., 2021). Building on the invariance principle, Invariant Causal Prediction formalizes causal variable selection via stability across environments (Peters et al., 2016). A rich line of identifiability results shows that with auxiliary variables or weak interventions, nonlinear ICA and VAE-based approaches can recover latent sources up to permissible transformations (Hyvarinen et al., 2019; Khemakhem et al., 2020). For vision, structured VAEs with causal priors explicitly separate content from style/context (Yang et al., 2021), and temporal–interventional setups such as Citris leverage intervention sequences to learn identifiable factors (Lippe et al., 2022). Beyond model proposals, empirical and theoretical analyses caution that purely unsupervised disentanglement is generally impossible without inductive biases or weak supervision, motivating the use of environment cues or metadata (Locatello et al., 2019). One line of research leverages distributional shifts in latent variables, where nonstationary or environmental variation facilitates the recovery of nonlinear independent components and the underlying causal structures (Kong et al., 2022; Zhang and Hyvärinen, 2011; Chen et al., 2024). Our approach follows this CRL perspective: we infer a latent space with a VAE, factorize it into invariant content and bias/context, then exploit the bias variables for environment-aware prediction while aligning them with invariant predictions.
Label shift in OOD problems Label shift occurs when the label prior $p(y)$ changes between training and test domains while the class-conditional $p(x\mid y)$ remains stable (Lipton et al., 2018). Classical corrections include EM-based prior adjustment (Saerens et al., 2002) and Black-Box Shift Estimation (BBSE), which uses a calibrated source classifier’s confusion matrix to estimate target priors and reweight predictions (Lipton et al., 2018). Subsequent work unifies moment-matching and maximum-likelihood estimators and emphasizes the central role of calibration (e.g., bias-corrected temperature scaling, BCTS) for reliable prior estimation (Garg et al., 2020; Azizzadenesheli et al., 2019). In streaming or nonstationary settings, online label-shift adaptation updates class-prior estimates on the fly without access to target labels (Wu et al., 2021; Bai et al., 2022). Benchmarks such as WILDS and Wild-Time document that label shift frequently co-occurs with input shift in real applications (e.g., medical imaging, wildlife monitoring, and temporally evolving text) (Koh et al., 2021; Yao et al., 2022). Consistent with this evidence, we explicitly account for changes in $p(y)$ when adapting to new environments, combining prior correction with environment-aware mixture-of-experts weighting informed by the biased latent variables.
Unsupervised domain adaptation Unlike DG, unsupervised domain adaptation utilizes unlabeled target data to guide model adaptation. Due to the observation of target distribution, many methods focus on distribution alignment by minimizing domain divergence (Rozantsev et al., 2018; Sun and Saenko, 2016; Kang et al., 2019; Liu et al., 2020), using importance reweighting (Jiang and Zhai, 2007; Xu et al., 2019; Fang et al., 2020), and learning with adversarial discriminators, such as DANN (Ganin et al., 2016), ADDA (Tzeng et al., 2017), and WDDA (Shen et al., 2018). Beyond distribution alignment, some methods focus on invariant information-based adaptation, including self-training (Lee and others, 2013; Zou et al., 2018; Saito et al., 2020; Chen et al., 2020; Kumar et al., 2020; Liu et al., 2021a), which uses pseudo-labels from a source-trained model to guide target domain learning, and batch normalization-based regularization (Li et al., 2018b; Liu et al., 2021b), which adapts only BN layers while keeping other parameters invariant. Recently, disentangled representation learning (Kong et al., 2022; Li et al., 2023b) has shown strong potential for adaptation by separating domain-invariant content from domain-specific bias.
Test-Time domain adaptation In real-world scenarios, the target domain distribution is often unknown in advance, making it challenging to train an adaptive model beforehand, which motivates test time adaptation. The typical solution is Test-Time Training (TTT) (Sun et al., 2020; Liu et al., 2021c; Mirza et al., 2023; Li et al., 2023a), which adapts models during inference by optimizing the same self-supervised learning objective used during pre-training. Besides, even without access to the target distribution, adaptation remains feasible with invariant information learned from the source domain, such as through self-training (Chen et al., 2022; Tomar et al., 2023; Su et al., 2022; Goyal et al., 2022) or network regularization (Wang et al., 2020a; Liang et al., 2020). The most closely related work to ours is Stable Feature Boosting (SFB), which leverages unstable features alongside invariant representations to generate pseudo-labels when only the image is affected by environmental factors. Our approach differs in three key ways. First, we delve deeper into latent variable modeling to provide a more comprehensive explanation of bias. Second, our model accommodates more complex bias scenarios, allowing the environment variable to influence both the image and the label. Third, we integrate invariant correction and expert reweighting to effectively utilize biased data.
Appendix B Proof and Discussion of section 2
B.1 Proof of the identification of latent variables
Lemma B.1.
@@ B.1 Proof of the identification of latent variables
  • A1 (Smooth and positive density): $p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})$ is smooth with strictly positive density on its support.
  • A2 (Conditional independence and content invariance): Given $(\mathbf{e},\mathbf{y})$, the coordinates of $\mathbf{z}$ are independent,
    $\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i};\mathbf{e},\mathbf{y}),$
    where $n=n_{b}+n_{c}$, $z_{1},\ldots,z_{n_{b}}$ correspond to $\mathbf{b}$ and $z_{n_{b}+1},\ldots,z_{n}$ correspond to $\mathbf{c}$. Moreover, content coordinates are environment-invariant: for $i>n_{b}$, $q_{i}(c_{i-n_{b}};\mathbf{e},\mathbf{y})=q_{i}(c_{i-n_{b}};\mathbf{y})$.
  • A3 (Linear independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$ (and the relevant $\mathbf{y}$), there exist $n_{b}+1$ environments $e_{0},e_{1},\ldots,e_{n_{b}}$ such that the $n_{b}$ vectors
    $\bm{w}(\mathbf{b},e_{j},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y}),\quad j=1,\ldots,n_{b},$
    are linearly independent, where
    $\bm{w}(\mathbf{b},e_{i},\mathbf{y}):=\Big(\tfrac{\partial q_{1}(b_{1};e_{i},\mathbf{y})}{\partial b_{1}},\ldots,\tfrac{\partial q_{n_{b}}(b_{n_{b}};e_{i},\mathbf{y})}{\partial b_{n_{b}}}\Big)^{\top}.$
  • A4 (Domain variability): There exist $e_{i}\neq e_{j}$ such that, for the relevant $\mathbf{y}$ and for any measurable set $A_{\mathbf{z}}\subseteq\mathcal{Z}$ with non-zero probability that cannot be written as $\Omega_{\mathbf{c}}\times\mathcal{B}$ for any $\Omega_{\mathbf{c}}\subset\mathcal{C}$,
    $\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{i},\mathbf{y})\,d\mathbf{z}\;\neq\;\int_{A_{\mathbf{z}}}p(\mathbf{z}\mid\mathbf{e}_{j},\mathbf{y})\,d\mathbf{z}.$
[…]
Proof.
We start from the matched conditional distribution $p_{\hat{\mathbf{x}}\mid\mathbf{e},\mathbf{y}}=p_{\mathbf{x}\mid\mathbf{e},\mathbf{y}}$ and introduce the diffeomorphism
$h:=g^{-1}\circ\hat{g}:\;\hat{\mathbf{z}}\mapsto\mathbf{z}.$
Then, for all $(\mathbf{e},\mathbf{y})$,
$p_{\hat{g}(\hat{\mathbf{z}})\mid\mathbf{e},\mathbf{y}}=p_{g(\mathbf{z})\mid\mathbf{e},\mathbf{y}}.$ (8)
By the change-of-variables formula,
@@
By A2 (conditional independence), letting $q_{i}(z_{i};\mathbf{e},\mathbf{y}):=\log p_{z_{i}\mid\mathbf{e},\mathbf{y}}(z_{i})$ and $\hat{q}_{i}(\hat{z}_{i};\mathbf{e},\mathbf{y}):=\log p_{\hat{z}_{i}\mid\mathbf{e},\mathbf{y}}(\hat{z}_{i})$, we have
$\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i};\mathbf{e},\mathbf{y}),\quad\log p_{\hat{\mathbf{z}}\mid\mathbf{e},\mathbf{y}}(\hat{\mathbf{z}}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i};\mathbf{e},\mathbf{y}).$ (10)
[…]
$\sum_{i=1}^{n}q_{i}\big(z_{i};\mathbf{e},\mathbf{y}\big)\;+\;\log\big|\det\mathbf{J}_{h}(\hat{\mathbf{z}})\big|\;=\;\sum_{i=1}^{n}\hat{q}_{i}\big(\hat{z}_{i};\mathbf{e},\mathbf{y}\big).$ (11)
Step 1: $\mathbf{B}=\partial\mathbf{b}/\partial\hat{\mathbf{c}}=\mathbf{0}$.
Differentiate equation 11 with respect to $\hat{c}_{j}$ for any $j\in\{1,\ldots,n_{c}\}$:
$\sum_{i=1}^{n}\frac{\partial q_{i}(z_{i};\mathbf{e},\mathbf{y})}{\partial z_{i}}\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}\;+\;\frac{\partial}{\partial\hat{c}_{j}}\log\big|\det\mathbf{J}_{h}(\hat{\mathbf{z}})\big|\;=\;\frac{\partial\hat{q}_{j}(\hat{c}_{j};\mathbf{e},\mathbf{y})}{\partial\hat{c}_{j}}.$ (12)
Note that $\mathbf{J}_{h}$ does not depend on $\mathbf{e}$, since $h=g^{-1}\circ\hat{g}$ is a mapping between latent spaces only. Fix $e_{0},\ldots,e_{n_{b}}$ as in A3 and subtract the version of equation 12 at $e_{0}$ from that at $e_{k}$ ($k=1,\ldots,n_{b}$); the Jacobian term cancels:
$\sum_{i=1}^{n}\!\left(\frac{\partial q_{i}(z_{i};e_{k},\mathbf{y})}{\partial z_{i}}-\frac{\partial q_{i}(z_{i};e_{0},\mathbf{y})}{\partial z_{i}}\right)\!\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}\;=\;\frac{\partial\hat{q}_{j}(\hat{c}_{j};e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j};e_{0},\mathbf{y})}{\partial\hat{c}_{j}}.$ (13)
By A2, for content coordinates ($i>n_{b}$) the difference on the left is zero because $q_{i}$ does not depend on $\mathbf{e}$. By the definition of the learned invariant block, $p_{\hat{c}_{j}\mid\mathbf{e},\mathbf{y}}$ is invariant in $\mathbf{e}$, hence the right-hand side of equation 13 is zero. Therefore,
$\sum_{i=1}^{n_{b}}\!\left(\frac{\partial q_{i}(b_{i};e_{k},\mathbf{y})}{\partial b_{i}}-\frac{\partial q_{i}(b_{i};e_{0},\mathbf{y})}{\partial b_{i}}\right)\!\cdot\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0,\quad k=1,\ldots,n_{b}.$
By A3, the $n_{b}$ vectors $\bm{w}(\mathbf{b},e_{k},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y})$ are linearly independent, so the only solution is
$\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0,\qquad\forall\,i\in\{1,\ldots,n_{b}\},\;j\in\{1,\ldots,n_{c}\}.$
Let the Jacobian of $h$ be block-partitioned as
$\mathbf{J}_{h}=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix}=\begin{bmatrix}\partial\mathbf{b}/\partial\hat{\mathbf{b}}&\partial\mathbf{b}/\partial\hat{\mathbf{c}}\\ \partial\mathbf{c}/\partial\hat{\mathbf{b}}&\partial\mathbf{c}/\partial\hat{\mathbf{c}}\end{bmatrix}$.
[…]
Step 2: $\mathbf{C}=\partial\mathbf{c}/\partial\hat{\mathbf{b}}=\mathbf{0}$.
[…]
B.2 Discussion of Lemma 2.1
Concepts Explanation
We would like to detailedly explain some concepts in Lemma 2.1 as follows: (1)Here “Positive support” means the set of all points where a probability distribution assigns strictly positive probability or density (i.e., $p(x)>0$). This is emphasized because the probability distribution may have some point in 0 for example the distribution are uniform distribution. To maintain academic rigor, this hypothesis was highlighted. (2)The $\mathcal{B}$ is introduced in A3 and it refers to the set (or space) of domain-specific latent variables. (3)Linear independence means that changes in different environments produce distinct, non-redundant effects on the domain-specific latent variable $b$; none of these variations can be expressed as a combination of others. This is a reasonable assumption because, with multiple diverse environments, it ensures that each domain provides unique information about how $b$ varies, making the latent factors identifiable rather than entangled. The detailed explanations of this lemma can be found in paper (Kong et al., 2022).
Discussion of Assumptions.
(A1) Smooth and positive density. This assumption requires that the conditional densities $p({\mathbf{z}}\mid e,y)$ are smooth and strictly positive on their supports. Intuitively, it ensures that the latent space is continuous and that every possible latent configuration has a non-zero probability of occurrence. In practice, this assumption is easily satisfied: with sufficient data, empirical or neural-network-based density estimations naturally produce smooth and positive densities. Since our framework adopts a VAE-based generative model, the latent distribution is modeled by Gaussian families whose densities are positive everywhere, satisfying this assumption automatically. However, this assumption may fail when the latent space is discrete or piecewise constant, such as in deterministic autoencoders or models that collapse the latent variance to zero. In those degenerate cases, $p({\mathbf{z}}\mid e,y)$ no longer has a smooth positive density, and the identifiability conditions in Lemma 2.1 no longer hold rigorously.
(A2) Conditional independence and content invariance. This assumption states that the coordinates of the latent variables ${\mathbf{z}}=(z_{1},\dots,z_{n})$ are conditionally independent given $e$ and $y$, i.e., $\log p({\mathbf{z}}\mid e,y)=\sum_{i}q_{i}(z_{i};e,y)$. Conceptually, this means that each latent dimension corresponds to an independent generative factor once the environment and label are specified. It also implies that content variables are invariant across environments—only the bias-related coordinates are affected by domain change. This assumption is reasonable in our framework because, when the generative process is disentangled and the noise level is small, the variation of $\mathbf{z}$} can indeed be attributed to independent factors conditioned on $e$ and $y$. Moreover, the VAE objective encourages conditional independence through the factorization of the approximate posterior. However, this assumption may be violated when latent variables interact strongly (e.g., through nonlinear dependencies) or when noise in the data introduces correlated variations between coordinates. In such cases, exact independence cannot be guaranteed, but approximate independence is often sufficient for stable learning and representation disentanglement.
(A3) Linear independence. This assumption imposes a technical requirement on the partial derivatives of the latent log-density functions with respect to environment changes. Intuitively, this means that the changes introduced by different environments are sufficiently diverse so that the bias components can be linearly separated from one another. In practice, this condition cannot be enforced directly, but it is often satisfied when multiple environments exist and differ in a sufficiently rich manner. Empirically, in synthetic datasets, when the number of environments is large, linear independence almost always holds, guaranteeing successful disentanglement. In real-world data, this condition may be partially violated because the number of distinct environments if we think this is domain number is limited. Nevertheless, prior works (Kong et al., 2022; Li et al., 2023b) and our empirical results show that even approximate linear independence is sufficient to achieve robust representations. This is may because the so-called domain in domain generalization are large domain and in this domain that there existing small domains to satisfy the assumptions.
(A4) Domain variability. This assumption requires that, for at least one pair of distinct environments $e_{i}\neq e_{j}$ and the corresponding class $y$, there exists a measurable subset $A_{z}\subset\mathcal{Z}$ with non-zero probability that cannot be expressed as a purely content-aligned region $\Omega_{c}\times B$, such that
$\int_{A_{z}}p({\mathbf{z}}\mid e_{i},y)\,dz\;\neq\;\int_{A_{z}}p({\mathbf{z}}\mid e_{j},y)\,dz.$
This ensures that the conditional latent distribution genuinely varies across domains, capturing environment-dependent factors beyond the content component. Because the mapping ${\mathbf{x}}=g({\mathbf{z}})$ is deterministic and noiseless, $p({\mathbf{x}}\mid e,y)$ is the push forward of $p({\mathbf{z}}\mid e,y)$; therefore, if $p({\mathbf{z}}\mid e_{i},y)=p({\mathbf{z}}\mid e_{j},y)$, we would also have $p({\mathbf{x}}\mid e_{i},y)=p({\mathbf{x}}\mid e_{j},y)$. However, this equality does not hold in domain generalization, where environments are explicitly assumed to induce observable shifts. To see this, consider that if the joint distributions were identical across environments,
$p({\mathbf{x}},y,e_{i})=p({\mathbf{x}},y,e_{j}),$
then by the chain rule,
$p({\mathbf{x}},y,e)=p({\mathbf{x}}\mid y,e)\,p(y\mid e)\,p(e).$
Integrating both sides over a measurable subset $A_{x}$ gives
$\int_{A_{x}}p({\mathbf{x}}\mid y,e_{i})\,p(y\mid e_{i})\,dx=\int_{A_{x}}p({\mathbf{x}}\mid y,e_{j})\,p(y\mid e_{j})\,dx.$
If $p({\mathbf{x}}\mid y,e_{i})=p({\mathbf{x}}\mid y,e_{j})$ held for all $A_{x}$, then equality would require $p(y\mid e_{i})=p(y\mid e_{j})$. Yet, in our setting, we explicitly consider label shift, meaning $p(y\mid e)$ varies with $e$; hence, this equality cannot generally hold. Consequently, A4 guarantees genuine domain variability at the level of $p({\mathbf{z}}\mid e,y)$, which propagates to $p({\mathbf{x}}\mid e,y)$ through the deterministic mapping. Nevertheless, A4 may fail under degenerate or unrealistic conditions, such as:
  1. (i)
    when environments do not induce any shift in latent factors ($p({\mathbf{z}}\mid e_{i},y)=p({\mathbf{z}}\mid e_{j},y)$ for all $i,j$), making domains statistically indistinguishable;
  2. (ii)
    when the mapping $g$ is stochastic or non-injective, collapsing different latent codes into identical observations, which hides latent variability in the observed space;
  3. (iii)
    or when label shift is absent ($p(y\mid e_{i})=p(y\mid e_{j})$), eliminating one key source of distributional difference across domains.
These cases violate the core assumption that domains differ in meaningful, non-content-related ways and thus fall outside our framework.
B.3 Better predictor with adding bias
Proof.
Under Assumptions A1–A4, we compare the Bayes-optimal risks for predictors based on $c$ versus $(c,b)$.
Log‐loss.
The Bayes-optimal log-loss risks satisfy
where
There is a positive-measure set of $(c,b)$ on which $p(y\mid c,b)\neq p(y\mid c)$. Hence
so $H(Y\mid C,B)<H(Y\mid C)$ and therefore
Squared‐error loss.
The Bayes‐optimal squared‐error risks satisfy
By the law of total variance,
Since $B$ influences the conditional mean of $Y$ given $C$, the second term is strictly positive on a set of positive probability. Thus
and consequently
This completes the proof that incorporating $b$ strictly reduces both log‐loss and squared‐error Bayes risks. ∎
B.4 The derivation of the KL optimization
We want to solve the following problem:
Let
By the definition of KL divergence, for discrete $y$:
Hence, for each $e$,
Therefore,
To find the minimizing $q$, we set
$\frac{\partial\mathcal{J}}{\partial q(y\mid{\mathbf{b}})}\;=\;0.$
First, note
Only the term $\log q(y\mid{\mathbf{b}})$ depends on $q$. For a specific $y$,
Hence, for each $e$:
Summing over $e$,
Also,
Hence,
Setting this to zero:
$\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda=0,$
which gives
Thus
We require $\sum_{y}q(y\mid{\mathbf{b}})=1$. Then:
Since $\sum_{y}p(y\mid{\mathbf{b}},e)=1$ for each $e$, we get
Hence
$1=\frac{1}{\lambda}\cdot 1\quad\Longrightarrow\quad\lambda=1.$
Therefore,
$q(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
B.5 Decomposition and invariant
$=\frac{p(\mathbf{c},\mathbf{b},y=1)}{p(\mathbf{c},\mathbf{b},y=0)}$
$\forall e_{i},e_{j}\in e_{all}$ (15)
$\quad\mathbf{c}\textbf{ only depends on }y$
${p(\mathbf{c}\mid y=1,e=e_{i})}=p(\mathbf{c}\mid y=1,e=e_{j})$
Use Bayes rule
B.6 Proof of Equation 1